OpenStudy (anonymous):

is any one good at transformations and coordinates in math? if you are can you help me

jagr2713 (jagr2713):

Well you can post it in the Math section and i am sure someone will help :D openstudy.com/study#/groups/Mathematics

OpenStudy (anonymous):

i did and no one has come to help me for a hour

OpenStudy (mle):

ask in math basic idea y = m(x - h) + k m is the slope h is how much left or right k is how much up or down that was the point slope form of a linear equation for a quadratic it is almost exactly the same y = a(x - h)^2 + k a is how much stretched h is how much left or right (again) k is how much up or down (again) notice a pattern yet? for equation of circle (x - h)^2 + (y - k)^2 = r^2 r is the radius h is how much left or right (again) k is how much up or down (again)

DivineSolar (divinesolar):

@blackdragon316 That does not mean post in other sections for a higher chance.